5 average office workers earn a total average monthly salary of $10,000.If the average salary of two of these workers total $13,000,what is the average monthly income of the remaining three workers?
Found 2 solutions by stanbon, sachi:
Answer by stanbon(75887)

5 average office workers earn a total average monthly salary of $10,000.If the average salary of two of these workers total $13,000,what is the average monthly income of the remaining three workers?
---
Total earnings by 5 workers = 5*10000 = 50,000
---
Total salary of 2 workers = 2*13000 = 26000
----
Average of 3 remaining workers = (50,000-26000)/3 = 24000/3 = $8000
=======================
Cheers,
Stan H.
=======================

Answer by sachi(548)

5 average office workers earn a total average monthly salary of $10,000.
so their total salary=5*10000=$50000
the average salary of two of these workers total $13,000,
so the total salary of two workers=2*13000=&26000
the total salary of the three remaining workers=50000-26000=$24000
so the average monthly income of the remaining three workers =2400/3=&8000
ans